OpenStudy (anonymous):

a store bought milk for $2.50 per half gallon and stored it in two refrigerators. During the night one refrigerator malfunctioned and ruined 13 half gallons. If the remaining milk is sold for $3.94 per half gallon, how many half gallons did the store buy if they made a profit of $89.90?

OpenStudy (anonymous):

Find profit per half gallon $3.94-$2.50 = $1.44 Find out how much money was lost because of fridge 13 * $2.50 = $32.50 Find out how much money the store needs to make to make $89.90 profit $32.50 + $89.90 = $122.4 Find how many half gallons bought $122.4/$1.44 = 85 85 Half gallons were bought by the store.

OpenStudy (anonymous):

let the number of half gallons bought = x cost price of x half gallons = 2.50x spoiled = 13 half gallons left = (x -13) half gallons selling price of (x - 13) gallons = (x-13)*3.94 = 3.94x - 51.22 profit = selling price - cost price = (3.94x - 51.22) - (2.50x) = 3.94x - 2.50x -51.22 = 1.44x - 51.22 ATQ, profit = 89.90 so 1.44x - 51.22 = 89.90 1.44x = 89.90 + 51.22 x = 141.12/1.44 x = 98 so 98 half-gallons were bought by the store

OpenStudy (anonymous):

checking the answer cost price of 98 half-gallons = 98*2.50 = 245 half-gallons sold = 98 -13 = 85 selling price of 85 half-gallons = 85*3.94 = 334.90 profit = selling price - cost price = 334.90 - 245 = 89.90